OpenStudy (tomektomeknyc):

For what values of theta belongs to (0,2) is the expression 1/ 1 + sec theta undefined? a) 2 pi/3, 4 pi/3 b) pi/2, 3pi/2 c) 0 d) 5 pi/6 , 7pi/6

OpenStudy (mathmale):

Hey, Tomcat, I suspect your interval is (0,2pi), and not (0,2). Double-check, would you please?

OpenStudy (tomektomeknyc):

It says (0, 2) but it might be wrong

OpenStudy (mathmale):

Here's the original expression:\[\frac{ 1 }{ 1+\sec \theta }\]

OpenStudy (mathmale):

This expression will be undefined for any theta for which the denominator, 1+sec theta, is zero, right? So, we must solve \[1+\sec \theta=0\]

OpenStudy (tomektomeknyc):

Sorry, I made a mistake, it supposed to be 1 / 2 + sec theta

OpenStudy (mathmale):

How would you go about solving this equation for theta, restricting theta to the interval (0,2pi)?

OpenStudy (mathmale):

OK: Then we must now solve \[2+\sec \theta=0\]

OpenStudy (mathmale):

How?

OpenStudy (tomektomeknyc):

sec theta = - 2, but it will not be in interval 0,2 pi

OpenStudy (mathmale):

TomKat: Here's what I'd do:\[2+\sec \theta=0\] \[\sec \theta=\frac{ -2 }{ 1 }\] \[\sec \theta=\frac{ 1 }{ \cos \theta }\]

OpenStudy (mathmale):

\[\frac{ -2 }{ 1 }=\sec \theta=\frac{ 1 }{ \cos \theta }\]

OpenStudy (mathmale):

Can you now solve this for theta? How many solutions would you expect, and why?

OpenStudy (mathmale):

Equivalently, \[\cos \theta=-1/2\]

OpenStudy (mathmale):

Solutions?

OpenStudy (tomektomeknyc):

theta = - 1/2 / cos

OpenStudy (mathmale):

TomKat: the cosine function is an operator (a command); it can't be separated from the angle (theta) in this manner. Review the definitions of the 3 basic trig functions. In particular, note that \[\cos \theta=\frac{ adjacent side }{ hypotenuse }=\frac{ adj }{ hyp}\]

OpenStudy (mathmale):

Does that look at all familiar to you?

OpenStudy (tomektomeknyc):

yes

OpenStudy (mathmale):

Good. Then perhaps it'd make more sense to write something like this: \[\cos \theta=\frac{ adj }{ hyp }=\frac{ -1 }{ 2 }\] Could you draw one or more triangles within a circle for which this relationship is true? In other words, choose the appropriate quadrants (there are 2 that apply here), and draw in each quadrant a triangle with adjacent side -1 and hypotenuse 2?

OpenStudy (mathmale):

If you're able to do that, it shouldn't be hard to identify the angles involved. We're trying to solve for the values of those two angles.

OpenStudy (tomektomeknyc):

that will apply to quadrant 2 and 3, thats where cos is negative

OpenStudy (tomektomeknyc):